# Two-level Nonlinear Preconditioning Methods for Flood Models Posed on Perforated Domains

### v1

@@ 6.3 Diffusive Wave Model on Large Realistic Urban Domain
In the final numerical experiment, referred to as Example 3, we consider the Diffusive Wave model (1) used to model a hypothetical flood in a densely urbanized area the city of Nice, France. The model domain is depicted in Figure 7 and has dimension $850\times 1500$ meters. In this test case scenario, the flood is produced by an overflow of Paillon river in the north-west part of the domain. Once again, the buildings are removed from the computational domain, leading to a large number of perforations (447 in total). The bathymetry $z_{b}$ uses the $1m$ Digital Elevation Model available from [18]. The parameters of the Diffusive Wave model are chosen formula as $\alpha=1.5$ and $\gamma=0.5$, based on Chézy’s formula, with friction coefficient $c_{f}=30$ which corresponding to a rough terrain [29]. We set Dirichlet boundary conditions, with water flow of about 2 metres coming from the Paillon river region and no water flow ($u=z_{b}$) elsewhere. The initial condition is taken as equal to $z_{b}$ and the final simulation time is set to $T_{f}=1500$ seconds, which corresponds to a steady state solution.
Consider the NRAS iteration (16), which contains the solution of local nonlinear subproblems via Newton’s method. In practice, for time-dependent problems, we come across the issue that some local subpbroblems will not converge for a given time increment $\Delta t_{n}=t_{n+1}-t_{n}$ in the residual function (5). We have found for our numerical examples that often, it is a small portion of subproblems that will not converge, generally due to stagnation in convergence. Therefore, it may be inefficient to reduce the global time increment for all subproblems and the entire iteration. We implement a local time step reduction strategy which makes it possible to reduce the time increment locally, then use this solution as an initial guess for the original system. This process is described in Algorithm 3.
@@
In this numerical experiment, with the exception of Newton’s method, the discussed local adaptive time-stepping method of Algorithm 3 is implemented with a global time increment of $\Delta t=10$ seconds for each time step for all methods. We find that with this local adaptive time stepping, a global time increment reduction is not necessary for convergence.
@@
The spatial discretization results in 81444 mesh points and 131581 triangles in the model domain. As mentioned, the fine-scale triangulation will be conforming to the polygonal coarse partitioning. Although we record results for various numbers of subdomains $N$, the fine-scale mesh will be kept consistent throughout the experiment and is conforming to a $N=8\times 16$ coarse mesh; therefore, it will conforming to $N=1\times 2$, $N=2\times 4$, $N=4\times 8$ meshes as well. The coarse and fine partitionings are shown in Figure 6 for multiple coarse partitionings. By the end of the simulation, 38216 points and 60846 triangles are “flooded” such that $h\geq 1$cm. The total number of coarse nodes are given by $N_{\cal V}=$795, 357, 126, and 34 for $N=8\times 16,4\times 8,2\times 4,$ and $1\times 2$ subdomains, respectively. The overlap is once again set to $\frac{1}{20}\mathcal{H}_{j}$.
@@
Figure 6: Coarse (thick blue lines) and fine (thin blue lines) discretization for $N=2\times 4$ (left) and $N=8\times 16$ (right) subdomains for model domain representing the Nice port and Paillon river.
@@
Recall that at every time step, the nonlinear system (5) has to be solved. For this example, in addition to Newton’s method given by Algorithm 1, we provide numerical results for the two-step method from Algorithm 4, the one-level RASPEN method from Algorithm 2, the two-level RASPEN method from Algorithm 3, and the Anderson acceleration method from Algorithm 5.
@@
Figure 8 displays the cumulative number of (successful) outer iterations over the simulation time, obtained for each of the methods for various numbers of subdomains $N$. In addition, Tables 8, 9, 10 and 11 report the total number of outer iterations and the number of successful time steps. In addition, whenever appropriate, Tables 8, 9, 10 and 11 reports the number of GMRES and NRAS iterations. Instead of GMRES iterations per outer iteration, we display the cumulative number of GMRES iterations summed over all time steps and all outer iterations. For the two-step and Newton methods, at each outer iteration, the linear system is solved using GMRES method combined with the linear two-level RAS preconditioner (14). The number of GMRES iterations per outer iteration averaged over each time step are reported at Figure 9. The overall (cumulative) number of GMRES iterations is reported in Tables 8, 9, 10, and 11 for $N=8\times 16,4\times 8,2\times 4,1\times 2$, respectively.
@@
Figure 7: Numerical solution of Example 3 at various time steps, where the numerical solution is presented by $\mathbf{h}=\mathbf{u}-z_{b}(\mathbf{x})$. Top: $t=10$ (left) seconds, $t=250$ (right) seconds. Bottom: $t=500$ (left) seconds, final $t=1500$ (right) seconds. “Flooded” region (where $\mathbf{h}\geq 1$cm) are shown in colour, with underlying bathymetry $z_{b}$ shown in the background in black and white.
@@
Figure 8: Example 3, cumulative outer iterations.Top: $N=1\times 2$ (left), $N=2\times 4$ (right). Bottom: $N=4\times 8$ (left), $N=8\times 16$ (right).
@@
Figure 9: Example 3, GMRES iterations at each global outer iteration over time. Top: $N=1\times 2$ (left), $N=2\times 4$ (right). Bottom: $N=4\times 8$ (left), $N=8\times 16$ (right).
Method T. steps Outer Its GMRES Its $N_{it,loc}$
@@
Anderson 150 2524 N/A 1.81
2-step 150 687 20029 1.51
1-level RASPEN 150 686 100453 1.53
2-level RASPEN 150 730 12117(+2907 coarse) 1.84
Table 8: Example 3, $N=8\times 16$ subdomains.
Method T. steps Outer Its GMRES Its $N_{it,loc}$
@@
Anderson 150 1902 N/A 2.44
2-step 150 638 13164 1.91
1-level RASPEN 150 624 33117 1.95
2-level RASPEN 150 650 8065(+2031 coarse) 2.39
Table 9: Example 3, $N=4\times 8$ subdomains.
Method T. steps Outer Its GMRES Its $N_{it,loc}$
@@
Anderson 150 1657 N/A 3.43
2-step 150 592 9483 2.72
1-level RASPEN 150 571 9040 2.89
2-level RASPEN 150 595 6328(+1755 coarse) 3.44
Table 10: Example 3, $N=2\times 4$ subdomains.
Method T. steps Outer Its GMRES Its $N_{it,loc}$
@@
Anderson 150 1225 N/A 3.89
2-step 150 597 7794 4.28
1-level RASPEN 150 477 3827 4.72
2-level RASPEN 150 515 3983(+1114 coarse) 4.79
Table 11: Example 3, $N=1\times 2$.
@@
From Figure 8, we see that the RASPEN variants (both 1 and two-level) as well as the two-step method produce the smallest number of outer iterations, and are very similar to one another in this regard. Additionally, we see that the Anderson acceleration changes the most as we increase $N$, with the outer iterations generally increasing. As expected, Newton’s method results in a higher number of outer iterations when compared to the RASPEN and two-step methods.
@@
We see from Figure 9 that as the number of subdomains is increased, the one-level RASPEN method is no longer scalable; in fact, the GMRES solves per iteration are significantly larger than the other methods such that they are out of the scope of the vertical axis in the figure (green dots in Figure 9). Additionally, as the number of subdomains grow, two-level RASPEN clearly becomes the method which contains the lowest average number of GMRES iterations per outer iteration. As far as Newton’s method preconditioned with the two-level linear RAS preconditioner (blue dots in Figure 9), the method is fairly scalable, as the number of GMRES iterations are not increasing proportionally with $N$. However, we do see a slight increase as the number of subdomains is increased. We remark additionally that the Two-step method produces very similar results when compared to the preconditioned Newton’s method, as the Two-step method also involves a linearly preconditioned system via the two-level linear RAS preconditioner.
@@
From Table 8, 9, 10, and 11, we see the results from Figures 8 and 9 summarized. Particularly, Newton’s method results in much larger numbers of outer iterations when compared to the nonlinearly preconditioned variants (with the exception of the Anderson acceleration as $N$ grows large). Additionally, the GMRES iterations in one-level RASPEN grow as $N$ grows. It appears that if as the number of subdomains grows, two-level RASPEN is the most efficient, produces the lowest number of GMRES iterations with a relatively small number of coarse linear solves. We remark again that the coarse linear solves are computationally less expensive a global fine-scale linear solve.
Overall, from our numerical experiments, we draw the following conclusions. As expected, one-level RASPEN is not a scalable method as the number of subdomains $N$ grows large; for this reason, we do not generally recommend its use. For a stationary problem, Newton’s method is much more sensitive to the initial guess than the other proposed methods, with iteration counts increasing by up to 50%. While Newton’s method preconditioned with the two-level RAS method produces a fairly scalable iteration, the addition of an NRAS solve in the two-step method greatly accelerates the convergence. Additionally, the two-step method is quite straightforward to implement, and in principle the global linear system can be done via a direct solve (provided the system is not too large). For this linearized system in the Two-step and Newton’s methods, the Trefftz space performs well as a component of the two-level RAS preconditioner. However, when implemented, the two-level RASPEN method further reduces the total number of GMRES iterations compared to the Two-step and Newton methods. Additionally, specifically for the porous media examples, the two-level RASPEN method provides a significant acceleration in terms of outer iterations. The Anderson acceleration of a coarse two-step method is also an interesting alternative to Newton’s method, reducing the total iteration count. However, we do not generally recommend it over the two-step and two-level RASPEN methods.
@@ 7 Summary and Future Work
This article has provided a numerical experiment of multiple nonlinear domain decomposition methods to solve nonlinear equations on perforated domains. Specifically, we have focused on a Diffusive Wave model on a domain with a large number of perforations; these perforations represent buildings and fall/fences in urban areas. We have found that nonlinear preconditioning techniques, particularly two-level RASPEN with the coarse space from [4], greatly reduce the number of outer iterations for both tested model problems. Not only are the number of outer iterations reduced when compared to Newton’s method, but the inner number of GMRES iterations per outer iteration are reduced as well. Additionally, the two-step method which adds a nonlinear RAS iteration to each Newton iteration, is straightforward to implement and greatly reduces the number of outer and GMRES iterations; in principle, the sparse large linear system in the Two-step method could be solved via a direct solve. Furthermore, we have found that local time step reduction is an efficient method to eliminate the need for a global time step reduction for all subdomains. Future work involves load-balancing of the complexity of the subdomains and parallel implementation to provide detailed runtimes.

### v2

@@ 6.3 Diffusive Wave Model on Large Realistic Urban Domain
In the final numerical experiment, referred to as Example 3, we consider the Diffusive Wave model (1) used to model a hypothetical flood in a densely urbanized area of the city of Nice, France. The model domain is depicted in Figure 8 and has dimension $850\times 1500$ meters. In this test case scenario, the flood is produced by an overflow of Paillon river in the north-west part of the domain. Once again, the buildings are removed from the computational domain, leading to numerous perforations (447 in total). The bathymetry $z_{b}$ uses the $1m$ Digital Elevation Model available from [20]. The parameters of the Diffusive Wave model are chosen formula as $\alpha=1.5$ and $\gamma=0.5$, based on Chézy’s formula, with friction coefficient $c_{f}=30$ which corresponding to a rough terrain [32]. We set Dirichlet boundary conditions, with water flow of about 2 meters coming from the Paillon river region and $u=z_{b}$ elsewhere; this boundary condition is shown in Figure 6. The initial condition is taken as equal to $z_{b}$ and the final simulation time is set to $T_{f}=1500$ seconds, which corresponds to a steady state solution.
Figure 6: Boundary condition (thick blue and red lines) and corresponding model domain for Example 3. Red corresponds to $u=21.5$ meters, blue corresponds to $u=z_{b}$. This results in around a 2-meter elevation for the portion of the boundary representing the Paillon river.
Similarly to Example 2, we ensure that the fine-scale triangulation is identical as the number of subdomains $N$ changes by generating a finer background mesh to be used for various $N$. We generate a fine-scale triangulation for $N=8\times 16$ subdomains, then use this background triangulation for $N=1\times 2,2\times 4,4\times 8$ and $8\times 16$ subdomain partitionings. Figure 7 shows the matching background triangulation for $N=2\times 4$ and $8\times 16$ subdomains, respectively.
Consider the NRAS iteration (16), which contains the solution of local nonlinear subproblems via Newton’s method. In practice, for time-dependent problems, we come across the issue that some local subproblems will not converge for a given time increment $\Delta t_{n}=t_{n+1}-t_{n}$ in the residual function (5). We have found for our numerical examples that often, it is a small portion of subproblems that will not converge, generally due to stagnation in convergence. Therefore, it may be inefficient to reduce the global time increment for all subproblems and the entire iteration. We implement a local time step reduction strategy which makes it possible to reduce the time increment locally, then use this solution as an initial guess for the original system. This process is described in Algorithm 3.
@@
In this numerical experiment, except for Newton’s method, the discussed local adaptive time-stepping method of Algorithm 3 is implemented with a global time increment of $\Delta t=10$ seconds for each time step for all methods. We find that with this local adaptive time stepping, a global time increment reduction is not necessary for convergence.
@@
The spatial discretization results in 81444 mesh points and 131581 triangles in the model domain. As mentioned, the fine-scale triangulation will be conforming to the polygonal coarse partitioning. Although we record results for various numbers of subdomains $N$, the fine-scale mesh will be kept consistent throughout the experiment and is conforming to a $N=8\times 16$ coarse mesh; therefore, it will be conforming to $N=1\times 2$, $N=2\times 4$, $N=4\times 8$ meshes as well. The coarse and fine partitionings are shown in Figure 7 for multiple coarse partitionings. By the end of the simulation, 38216 points and 60846 triangles are “flooded” such that $h\geq 1$cm. The total number of coarse nodes are given by $N_{\cal V}=$795, 357, 126, and 34 for $N=8\times 16,4\times 8,2\times 4,$ and $1\times 2$ subdomains, respectively. The overlap is once again set to $\frac{1}{20}\mathcal{H}_{j}$.
@@
Figure 7: Coarse (thick blue lines) and fine (thin blue lines) discretizations for $N=2\times 4$ (left) and $N=8\times 16$ (right) subdomains for model domain representing the Nice port and Paillon river.
@@
Recall that at every time step, the nonlinear system (5) has to be solved. For this example, in addition to Newton’s method given by Algorithm 1, we provide numerical results for the One-level RASPEN method from Algorithm 2, the Two-level RASPEN method from Algorithm 3, the Two-step method from Algorithm 4, and the Anderson acceleration method from Algorithm 5.
Figure 9 displays the cumulative number of (successful) outer iterations over the simulation time, obtained for each of the methods for various numbers of subdomains $N$.
@@
In addition, we record various values in Tables 8, 9, 10 and 11. For this time-dependent example, these values report the total number of time steps $N_{T}$, the total number of outer iterations over time $N_{T}\cdot N_{k}$, and, whenever appropriate, also report the cumulative number of GMRES iterations over time $N_{T}\cdot N_{k}\cdot N_{\text{GM}}$ and the total number of local linear solves on a given subdomain over all time steps and outer iterations $N_{T}\cdot N_{k}\cdot N_{it,loc}$. For the Two-step and Newton methods, at each outer iteration, the linear system is solved using GMRES method combined with the linear two-level RAS preconditioner (14).
Additionally, we report the number of GMRES iterations per outer iteration $N_{\text{GM}}$ for each time step in Figure 10. As mentioned, the overall (cumulative) number of GMRES iterations is reported in Tables 8, 9, 10, and 11 for $N=8\times 16,4\times 8,2\times 4,1\times 2$, respectively.
@@
Figure 8: Numerical solution of Example 3 at various time steps, where the numerical solution is presented by $\mathbf{h}=\mathbf{u}-z_{b}(\mathbf{x})$. Top: $t=10$ (left) seconds, $t=250$ (right) seconds. Bottom: $t=500$ (left) seconds, final $t=1500$ (right) seconds. “Flooded” region (where $\mathbf{h}\geq 1$cm) are shown in color, with underlying bathymetry $z_{b}$ shown in the background in black and white.
@@
Figure 9: Example 3, cumulative outer iterations over time. Top: $N=1\times 2$ (left), $N=2\times 4$ (right). Bottom: $N=4\times 8$ (left), $N=8\times 16$ (right).
@@
Figure 10: Example 3, average number of GMRES iterations per outer iteration $N_{\text{GM}}$ at each time step. Top: $N=1\times 2$ (left), $N=2\times 4$ (right). Bottom: $N=4\times 8$ (left), $N=8\times 16$ (right).
Method $N_{T}$ $N_{T}\cdot N_{k}$ $(N_{T}\cdot N_{k}\cdot N_{\text{GM}})$ $N_{T}\cdot N_{k}\cdot N_{it,loc}$
@@
Anderson 150 2524 N/A 6849
2-step 150 687 20029 1557
1-level RASPEN 150 686 100453 1574
2-level RASPEN 150 730 12117(+2907) 2017
Table 8: Example 3, $N=8\times 16$ subdomains. The value in brackets represents $N_{t}\cdot N_{k}\cdot N_{c}$.
Method $N_{T}$ $N_{T}\cdot N_{k}$ $(N_{T}\cdot N_{k}\cdot N_{\text{GM}})$ $N_{T}\cdot N_{k}\cdot N_{it,loc}$
@@
Anderson 150 1902 N/A 6967
2-step 150 638 13164 1827
1-level RASPEN 150 624 33117 1824
2-level RASPEN 150 650 8065(+2031) 2334
Table 9: Example 3, $N=4\times 8$ subdomains. The value in brackets represents $N_{t}\cdot N_{k}\cdot N_{c}$.
Method $N_{T}$ $N_{T}\cdot N_{k}$ $(N_{T}\cdot N_{k}\cdot N_{\text{GM}})$ $N_{T}\cdot N_{k}\cdot N_{it,loc}$
@@
Anderson 150 1657 N/A 8538
2-step 150 592 9483 2412
1-level RASPEN 150 571 9040 2478
2-level RASPEN 150 595 6328(+1755) 3066
Table 10: Example 3, $N=2\times 4$ subdomains. The value in brackets represents $N_{t}\cdot N_{k}\cdot N_{c}$.
Method $N_{T}$ $N_{T}\cdot N_{k}$ $(N_{T}\cdot N_{k}\cdot N_{\text{GM}})$ $N_{T}\cdot N_{k}\cdot N_{it,loc}$
@@
Anderson 150 1225 N/A 7132
2-step 150 597 7794 3831
1-level RASPEN 150 477 3827 3377
2-level RASPEN 150 515 3983(+1114) 3698
Table 11: Example 3, $N=1\times 2$ subdomains. The value in brackets represents $N_{t}\cdot N_{k}\cdot N_{c}$.
@@
From Figure 9, we see that the RASPEN variants (both one and two-level) as well as the Two-step method produce the smallest number of total outer iterations, and are very similar to one another in this regard. Additionally, we see that the Anderson acceleration changes the most as we increase $N$, with the outer iterations generally increasing. As expected, Newton’s method results in a higher number of outer iterations when compared to the RASPEN and Two-step methods.
@@
We see from Figure 10 that as expected, as the number of subdomains is increased, the One-level RASPEN method is no longer scalable in terms of GMRES iterations; in fact, as $N$ grows, the average GMRES iterations $N_{\text{GM}}$ are significantly larger than the other methods such that they are out of the scope of the vertical axis in the figure (green dots in Figure 10). Additionally, as the number of subdomains grow, the Two-level RASPEN method clearly becomes the method with the lowest value $N_{\text{GM}}$ per outer iteration over time. As in the previous examples, for Newton’s method preconditioned with the two-level linear RAS preconditioner (blue dots in Figure 10) and the Two-step method preconditioned with the two-level linear RAS preconditioner (orange dots in Figure 10), the number of GMRES iterations are not increasing proportionally with $N$. However, we do see a slight increase as the number of subdomains is increased. We additionally remark that over time, the average number of GMRES iterations per outer iteration $N_{\text{GM}}$ in Figure 10 increases for each method besides the Two-level RASPEN method; this is generally most likely due to an increased proportion of the model domain being “flooded”. That is, the problem becomes increasingly difficult over time. However, this also suggests that the performance of the coarse correction is improved when applied “nonlinearly”, as is done in the Two-level RASPEN method.
@@
From Table 8, 9, 10, and 11, we see the results from Figures 9 and 10 summarized. Particularly, Newton’s method results in much larger numbers of outer iterations when compared to the nonlinearly preconditioned variants (with the exception of the Anderson acceleration as $N$ grows large). Additionally, the total number of GMRES iterations $N_{T}\cdot N_{k}\cdot N_{\text{GM}}$ in the One-level RASPEN method grows significantly as $N$ is increased, more than doubling as $N$ is doubled in one direction. It appears that particularly as the number of subdomains grows, Two-level RASPEN is the most efficient, producing the lowest number of total GMRES iterations with a relatively small number of coarse linear solves. We note again that the coarse linear solves are computationally less expensive than a global fine-scale linear solve. However, we do remark that we do not see the same type of acceleration in terms of outer iterations for Two-level RASPEN like we obtained in the other two numerical examples. Overall, particularly if one is choosing a high value of $N$, it appears that the Two-level RASPEN method outperforms the others.
We have mentioned that from Figure 9, the outer iterations of the Anderson method appear to have noticeably more dependence on $N$ than the outer iterations of the other methods. To view this phenomenon, we provide the average outer iterations per time step $N_{k}$ for Anderson’s method in Figure 11. We see from Figure 11 that the outer iterations do increase noticeably with $N$. But in fact, the number of outer iterations stays reasonably stable as time increases, and the pattern of outer iterations seems to match that of average GMRES iterations of two-level RASPEN in Figure 10. This dependence on $N$ suggests that the coarse space could be improved for this model problem.
Figure 11: Example 3, Anderson outer iterations per time step. Colors correspond to various coarse partitionings.
@@ 7 Summary and Future Work
This article has provided multiple efficient nonlinear domain decomposition methods to solve nonlinear equations on complex, perforated domains. Specifically, we have focused on a Diffusive Wave model on a domain with numerous perforations; these perforations represent buildings and walls/fences in urban areas. The coarse space form [4], designed for linear problems on perforated domains, proves to form efficient two-level methods for our difficult model problem.
From our numerical experiments, we draw the following conclusions. In terms of the number of GMRES iterations, as expected, the One-level RASPEN method is not a robust method as the number of subdomains $N$ grows large. For the linearized systems resulting at each iteration of the Two-step and Newton methods, the Trefftz space performs well as a component of the two-level RAS preconditioner and results in a reasonable number of inner GMRES iterations per outer iteration. However, these iterations slightly increase as $N$ is increased. This effect could be mitigated by updating the basis functions in the Trefftz space as the linearized system is updated. The Two-level RASPEN method further reduces the total number of GMRES iterations compared to the Two-step method, particularly as $N$ grows. However, in terms of total GMRES iterations $N_{k}\cdot N_{\text{GM}}$, due to the large number of $N_{k}$ for Newton’s method, Newton’s method results overall in a much larger total number of GMRES iterations.
In terms of total outer iterations, Newton’s method is much more sensitive to the initial guess than the other proposed methods in our experiment, with iteration counts varying by up to 50%. Additionally, Newton’s method generally results in extremely large iteration counts when compared to the Two-step and RASPEN variants. The addition of an NRAS solve in the Two-step method greatly accelerates convergence in terms of outer iterations when compared to Newton’s method. Specifically, for the stationary examples, the Two-level RASPEN method provides a significant acceleration in terms of outer iterations when compared to the other proposed methods. We do not see the same effect in the time-dependent example, likely due to the time stepping process and choice of the time increment. The Anderson acceleration of a coarse Two-step method is also an interesting alternative to Newton’s method, reducing the total iteration count, particularly for the stationary examples. However, for the time-dependent example, we do see a dependence on $N$ in terms of outer iterations for this method. Each iteration of Anderson’s method involves a small least-squares problem and is cheaper than a Newton iteration. However, it does not appear to accelerate convergence when compared to methods such as the Two-level RASPEN method.
In terms of implementation, both Newton’s method and the Two-step method involve a sparse linear system at each iteration which, in principle, could be solved via a direct solve (provided the system is not too large). Additionally, the Two-step method is quite straightforward to implement when compared to the Two-level RASPEN method, with the only addition from Newton’s method being a nonlinear RAS computation at each iteration. The Two-level RASPEN method involves a nonlinear RAS computation, a coarse nonlinear problem, and a global linear system which generally should be solved via a Krylov method. However, once implemented, the Two-level RASPEN method forms an extremely robust and efficient algorithm. Additionally, Anderson acceleration benefits from readily available built-in functions provided by multiple popular programming languages.
Finally, we have found that local time step reduction is an efficient method to eliminate the need for a global time step reduction for all subdomains. Future work involves a full parallel implementation to provide detailed CPU times.
Acknowledgments
This work has been supported by ANR Project Top-up (ANR-20-CE46-0005). The high-resolution structural data has been provided by Métropole Nice Côte d’Azur. We warmly thank Florient Largeron, chief of MNCA’s SIG 3D project, for his help in preparation of the data and for the multiple fruitful discussions.